# Language models should be subject to repeatable, open, domain-contextualized hallucination benchmarking

### v1

@@ 4 Case Study in Hallucination Dataset Development
We first develop sixty-four questions related to political procedures and medical diagnoses. For the political domain, we collaborated with the political practitioner to craft both the questions (e.g., “Can I register to vote on election day in California?” and “Do I need ID to vote in Arizona?”) and the corresponding factual, verifiable answers about electoral procedures (see e.g., ncsl2025election, ). In the medical domain, our research team generated questions (e.g., “I have ulcerative colitis. Will mesalamine help control it?” and “I’m experiencing joint pain, morning stiffness, and swollen fingers. Do I have psoriatic arthritis?”) using publicly available resources, and a physician provided plausible diagnoses.
@@
To generate LLM responses, we pass the sixty-four prompts to a single-level GPT-3.5 Turbo model via API.
@@
All responses are individually provided to a local (offline) keep performance indicator evaluation agent. This agent is capable of both reference-based hallucination measurement through traditional NLP metric computations, as well as reference-free hallucination measurement (leveraging a custom fine-tuned version of Llama 3.1 70B). This agent produces a slate of 12 scores per response. We describe these metrics and cite to the grounding source in Appendix A. For reach domain, we compute an aggregate score for each metrics, which is provided in Table 1.
5 Benchmark Results
@@
CTC RCR ROUGE-L BLEU METEOR THS
@@
CTC RCR ROUGE-L BLEU METEOR THS
@@
In the political example, the Factual Claim Density (FCD) score is higher for the LLM than for the GT, and the Factual Grounding References (FGR) is lower in the LLM compared to the GT. The interpretation is that more statements could be construed of as real facts (FCD) in the LLM compared to the GT, and that the LLM language appears to be less grounded in evidence (FGR) compared to the GT. In the medical example, the Factual Claim Density (FCD) score is lower for the LLM than for the GT, and the Factual Grounding References (FGR) is higher in the LLM compared to the GT. The interpretation is that fewer statements could be construed of as a real facts (FCD) in the LLM compared to the GT, and that the LLM language may appear to be more grounded in evidence compared to the GT (FGR).
Our assessment is that the physician’s responses, reflective of how he communicates with patients in real life, did not cite sources or explicitly reference underlying medical theory. By comparison, the LLM responses often grounded their claims in recognizable medical knowledge, referencing specific conditions, medications, or treatment protocols. When we followed up with the physician about this, he explained that doctors do not cite their sources in conversations with patients. This approach contrasts with how a political practitioner might cite evidence or policy when addressing questions about procedures (see, e.g. ncsl2025election, ). When we pressed the physician to think about what sources he would cite if he had to, he replied: “med school.” When pushed further, the physician explained that justifications for diagnoses are taught and memorized in medical school, then internalized over years of clinical practice.
@@
Our findings underscore the importance of aligning evaluation metrics with the specific task at hand. Several metrics — Fictional Disclaimer Frequency (FDF), Explicit Contextualization Score (ECS), Named Reference Ratio (NRR), Contextual Trigger Count (CTC) — scored $0.00.0$ for both GT and LLMs in both domains. This is particularly significant: it implies that neither the GT responses nor the LLM generations employed structured disclaimers, explicit contextualization, or named references. This absence is not likely attributable to model hallucination; instead it is potentially due to the nature of the measurement task itself. The sample prompts and reference responses often consist of straightforward factual statements that lack the stylistic or rhetorical cues these metrics are designed to detect. These bottom-coded hallucination metrics highlight a fundamental limitation in applying conventional fictionality-sensitive metrics to question and answer domains that are firmly grounded in fact rather than fiction.
@@
The text generated by the LLM mimics language that reflects greater realism or plausibility, but also includes language that is speculative. In the political example, the LLM responses averaged higher on the Real-world Concept Ratio (RCR) than the GT. This suggests that the LLM invoked more real-world concepts than the GT. However, this may not correspond to accurate grounding. The Speculative Claim Density (SCD) is higher for the LLM responses than for the GT, indicating an increased tendency to include language that hedges or speculates. The LLM produces text that has a higher THS than GT, driven in large part by the relatively elevated FCD and fewer contextual safeguards compared to the GT. The GT responses, in contrast, are shorter, more constrained, and structurally conservative — resulting in a low total hallucination score (THS), but also exhibit limited contextual sophistication (FGR, FDF, and ECS).
The board-certified, human-doctor produced responses that score considerably higher on THS than the model-generated text, principally due to the FCD chasm. A naive interpretation of THS would conclude that the doctor’s responses presented a greater risk of being hallucinated than the model-generated text. Obviously, this cannot be true. In fact, the way that doctors think, reason, and communicate about diagnoses from chief-complaints is simply poorly measured by these hallucination metrics. As doctors professionalize and develop medical expertise they develop heuristics that allow them to produce correct diagnoses quickly. Given limited, and fallible cognition, they prioritize getting answers correct, rather than citing the grounding scientific study that connects a complaint to a diagnoses. And, when faced with chief-complaints, doctors hold in mind a series of possible diagnoses — leading to higher values of SCD — some of which would be ruled out pending the result of a follow-up laboratory tests. For example, when responding to patients’ chief complaints and symptoms, doctors often say, “It’s possible you have some diagnosis. Let’s run some tests.”
@@
Consistent with our expectations, LLM outputs appear fluent, but they often deviate in surface form and phrasing from the human-generated reference text. Such findings underscore the limitations of current LLM generations in matching not just factual content, but precise expression and structure. Across both the political and medical domains, metrics that evaluate similarity between generated text and what a human would have written (i.e. ROUGE-L, BLEU and METEOR), GT responses score as near perfect matches to human written responses. In contrast, LLM responses receive considerably lower scores; there is substantial lexical and structural differences between model-generated text and human-written, ground-truth responses.
6 Alternative Positions
Alternative Position 1: Benchmarking is misguided.
Perhaps the negative consequences of hallucination errors are already so limited that the time and carbon spent reducing them could be better directed toward higher-impact problems. Or, perhaps since hallucinations can never be fully eliminated (li2023haluevallargescalehallucinationevaluation, ; banerjee2024llmshallucinateneedlive, ; xu2025, ), continued efforts to minimize them may represent a misallocation of resources. Another possibility is that general-purpose benchmarks will be sufficient to constrain models in specific tasks-settings, thereby minimizing hallucination.
None of these justifications seem tenable. Rising hallucination rates in current-generation, “reasoning” models contradict the notion that this is a moribund problem (Zhou2024larger_and_more, ); leaving a problem unaddressed because it is challenging or has unknown limit properties is inconsistent with the principles in computer science. These models are designed to produce output that is naturally legible to human users who apply them to both increasingly specialized, sophisticated tasks and broad, general tasks. We can be quite certain that the seemingly coherent language produced for general tasks will result in reduced scrutiny of all model-generated text (bender2021dangers, ). The decreased oversight of model output means that when a model hallucinates an incorrect answer, users often will not realize it: either becuase they lack the expertise to assess the response or because they do not devote sufficient cognitive effort to scrutinize it. If language models are used as “discount experts,” then the deleterious consequences of hallucinations are likely to be concentrated among vulnerable, resource-constrained populations. In light of these risks, expert-informed benchmarking remains a critical component of responsible language model development.
Alternative Position 2: Expertise is too expensive.
Perhaps the costs of benchmarking using domain-expert assessments outweigh the benefits of reducing hallucinations. Certainly, expertise is expensive — spot contracts for expert medical, legal and scientific consulting are commonly several hundreds of dollars per hour. What is more, expertise is highly domain specific. Gathering assessment and benchmarking tasks from experts in one domain is unlikely to prove useful even in closely-related, cognate domains. The implication, in the language of the drunkard’s search, is that it might be prohibitively expensive to generate enough lampposts of domain expertise to provide benchmark illumination across all areas of human knowledge.
We acknowledge these costs, but point out that they are minuscule compared to the enormous costs of training current-generation models and paying staff data scientists and engineers. Furthermore, we point out that model training produces models weights that are obsolescing, private goods; model weights are not commonly shared between companies or research teams, and current generation models are inevitably replaced by future generation models. In contrast, domain-expertise is nonrivalrous — the usage by one team does not decease the ability to be used by another team — and it is evergreen. Expertise can be reused to benchmark all generations of models, and is replaced only when the underlying science, law, or art evolves. The implication is that costs to gather expertise can be amortized across a relatively-longer time horizon and if this expertise is shared within an open scientific community and spread across many research groups, it can serve as a public good driving forward all research teams.
Alternative Position 3: Market dynamics will solve this problem.
The creation and provision of open, public data for domain-specific benchmarking is in contrast to closed, competitive market dynamics where firms compete to develop ever more performant language models. Perhaps the closed, competitive landscape is better suited to investing in benchmarks than the open scientific landscape.
@@ 6 Alternative Positions
While it is true that commercially developed language models are receiving considerable coverage in the press, it is important to remember that the theoretical foundations for these models were laid in open, scientific dialogue (e.g., vaswani2017attention, ). Commercial firms’ advantage has been in marshaling resources to glean text, process through compute to create private model weights, and deliver packaged product to consumers. Benchmarking is quite different from this: there are strong advantages to coordinating rather than competing, benchmarks are non-rivalrous and work best when commonly used within and across teams, and benchmark contains no trappings of being customer-facing and thus does not require a polished front-end. For exactly these reasons it is likely that market dynamics will lead to an under-provision of open domain-contextualized benchmark data, since this data could be used by a competitor to improve their models.
@@
There are some reasons to believe that dynamics that have led to the success of the open-source software paradigm may be applicable to expert benchmarking. Companies’ underlying value arises from applications built on top of core tools, and coordination around a common tool set (or benchmark set) creates the ability for collaboration both within and between firms in a way that supports and enhances development without giving away trade secrets. Like the open source software movement, companies that participate in open benchmarking benefit because they can recruit academic talent that has extensive exposure to tools and frameworks, rather than requiring costly in-house training on proprietary tooling. And, rather than the maintenance of benchmarks being served by a single team at a single organization, this maintenance can be supported by interested domain experts from the global community.
7 Conclusion
@@
In this position paper, we argue that subjecting language models to repeatable, open, and domain-contextualized hallucination benchmarks will advance language model development. Repeatability is essential for internal testing and scientific rigor because tuning models to perform well on narrow external benchmarks often leads to overfitting: strong performance on the benchmark task, but poor generalization to other contexts. Openness in benchmark development is foundational to science and enables users to independently verify claims made by software providers. Moreover, open benchmarks foster collaboration across organizations and between academic and industry groups, facilitate cost-sharing when procuring questions and answers, and invite contributions to be volunteered from domain experts, ensuring benchmarks reflect up-to-date knowledge. Domain-contextualized benchmarks that represent expert thinking are necessary to hallucination metrics remain effective and relevant when models are inevitably applied to increasingly challenging problems. Failure to include expertise risks deploying unhelpful, potential harmful language models.
Appendix A Metric Definitions
We report the Total Hallucination Score (THS) which is a composite measure that includes Factual Claim Density (FCD), Factual Grounding References (FGR), Fictional Disclaimer Frequency (FDF), Explicit Contextualization Score (ECS) [10]. FCD increases the possibility that a statement could contain a hallucination; grounding statements with a fact (FGR), or making a disclaimer that a statement is fictional (FDF & ECS) decreases the possibility that a statement contains a hallucination.
We also report conventional hallucination metrics: Speculative Claim Density (SCD), Named Reference Ratio (NRR), Contextual Trigger Count (CTC), and Real-world Concept Ratio (RCR).
Finally, we report machine summarization measures: Recall-Oriented Understudy for Gisting Evaluation – Longest Common Subsequence (ROUGE-L), Bilingual Evaluation Understudy (BLEU), and Metric for Evaluation of Translation with Explicit Ordering (METEOR). See Table 2 for a detailed description of the reported metrics.
@@ Appendix A Metric Definitions
FCD Factual Claim Density Number of claims that appear to be historical, scientific, or verifiable facts per 100 words. A lower FCD suggests fewer statements that could be mistaken for real facts.
@@
FGR Factual Grounding References Number of references to real-world evidence Lower FGR texts are more overtly fictional.
@@
FDF Fictional Disclaimer Frequency Frequency (per 100 tokens) of explicit indications the text is fictional. Higher FDF means texts are framed as not-based-in-fact.
@@
ECS Explicit Contextualization Score Frequency (per 100 tokens) text includes language such as “purely fictional” or “mythical.” Higher ECS indicates stronger distancing from factual framing.
@@
THS Total Hallucination Score Composite: FCD - (FGR + FDF + ECS). Higher THS implies a denser presence of fact-like claims that are not offset by disclaimers or real-world grounding.
@@
SCD Speculative Claim Density Proportion of sentences containing hedging or speculative language (e.g., “might,” “some believe,” “possibly”). Higher SCD may reflect caution in LLMs, but also signals epistemic uncertainty that can blur fact vs. fiction boundaries.
@@
NRR Named Reference Ratio Proportion of sentences that include named entities tied to identifiable sources (e.g., people, organizations, documents). Higher NRR suggests stronger citation patterns; low values often signal decontextualized or imagined references.
@@
CTC Contextual Trigger Count Number of explicit framing cues per 100 tokens (e.g., “according to legend,” “myth says,” “reportedly”). Higher CTC indicates the model is flagging uncertainty or fictionality more clearly.
@@
RCR Real-world Concept Ratio Proportion of noun phrases or terms that refer to concrete, verifiable real-world entities. Higher RCR reflects greater realism or plausibility; low values suggest fictional abstraction or vagueness.
@@
ROUGE-L Recall-Oriented Understudy for Gisting Evaluation Measures the longest common subsequence between generated and reference text. Higher ROUGE-L indicates stronger structural and lexical alignment with the reference.
@@
BLEU Bilingual Evaluation Understudy Precision-oriented score capturing overlap between n-grams in the candidate and reference. High BLEU reflects lexical fidelity to the reference; low BLEU suggests phrasing divergence.
@@
METEOR Metric for Evaluation of Translation with Explicit ORdering Harmonic mean of unigram precision and recall, with penalties for word order and semantic mismatch. More tolerant than BLEU; higher METEOR values indicate better fluency and synonym-aware alignment with the reference.
@@
Table 2: FCD, FGR, FDF, ECS, and the composite THS are drawn from Gosmar & Dahl, 2025 [10]. Additional reference-free metrics (SCD, NRR, CTC, RCR) [33, 55, 60, 45] extend interpretability in contexts lacking ground-truth. ROUGE-L [32], BLEU [44], and METEOR [2] are reference-based metrics standard in natural language evaluation.
Appendix B Performance Benchmarking
Determining how well an LLM performs is a complex problem. Very few industry or research leaders agree on the most important combination of measures or how they should be weighted. Some of the different ways to conceptualize performance include: time-to-train speed, time to n number of tokens predicted for inference or model efficiency measures such as the number of GPUs and memory size required to load and run a model successfully are all different ways of thinking about LLM performance [14]. In the paragraphs that follow, we describe the current state of the art approaches for performance benchmarking.
Training Performance Benchmarking
Due to the enormous compute memory and storage requirements some of the common resource measurements for establishing benchmarks of LLM training performance include: number of GPUs, size of memory, carbon footprint, and cost associated with electricity and personnel. The length of LLM training is primarily governed by the most efficient size of the token corpora to be processed, before overfitting occurs [14]. Interestingly, determining this optimal token size is done by testing training checkpoints on an accuracy metric. Of course, fine-tuning, mixture-of-experts (MoE) and other post-foundational training practices may extend the training process. From a practical sense, benchmarking test suites such as MLPerf provide tools for independent evaluation of LLM training processes[36].
Inference Performance Benchmarking
Some of the same performance constraints from training are applicable to inference — such as model size, memory, and GPU compute requirements. In addition to this, delivering an LLM’s output to a user (or more practically millions of concurrent users), in an acceptable time is both an engineering problem and a product issue. As such most performance benchmarks are highly subjective and are always local to the model architecture and hardware used. In order to achieve acceptable performance, trained models are often quantized and curated to either reduce the number of unused layers in the model and compress the model, or by reducing the precision of model weights [14, 61]. [^5] MLPerf offers well-documented benchmark tools for evaluation of LLM inference[47].
Appendix C Accuracy Benchmarking
Accuracy benchmarking is an actively developing research area [14, 29]. LLM systems vary in size, domain, technique, technology, applied use case and interface; as a result, evaluating the accuracy of model for accuracy is difficult, and no agreed-up measure or set of metrics exits. In the paragraphs that follow, we describe the current state of the art approaches for accuracy benchmarking.
Training Accuracy Benchmarking
Training accuracy depends primarily on the number of tokens, number of model parameters (model size) and number of GPU’s required to successfully complete the training process [14]. Rather than a single training run, frequently, model checkpoints are created after specific stopping criteria are met; then, metric snapshots from those checkpoints are used to determine future tuning and model optimization practices. Training performance is incredibly important to the field of GenAI and training outcomes are central to the in-the-wild performance of LLMs.
Inference Benchmarking
Hodak et al. highlight the emerging use of Recall-Oriented Understudy for Gisting Evaluation (ROGUE) [14] for text summarization task evaluation. ROGUE is a measure of similarity between human generated (gold standard) summarizations and the synthetic model outputs. Machine translation tasks are often evaluated using Bilingual Evaluation Understudy (BLEU) metrics and datasets, which also measures the similarity between reference translations and AI-generated output. In a more general sense, many research and industry applications leverage the 57 curated tasks from the Massive Multitask Language Understanding (MMLU) project as way to get a benchmark for LLM performance on common actions and knowledge [12]. In contrast, Zhou et al. actually introduce five time-of-inference benchmarks (addition, anagram, locality, science and transforms) which evaluate simple numeracy, vocabulary reshuffling, geographic knowledge, low-complexity to high-complexity scientific knowledge and multitask ability [61]. They then categorize those tasks with relationship to a given model’s ability answer correctly or avoid giving a direct answer. They they then implement these metrics and establish a set of benchmarks for each, based on human-judged difficulty of questions (prompts) given to a variety of models.

### v2

@@ 4.  Case Study in Hallucination Dataset Development
We first develop sixty-four questions related to political procedures and medical diagnoses. For the political domain, we collaborated with the political practitioner to craft both the questions (e.g., “Can I register to vote on election day in California?” and “Do I need ID to vote in Arizona?”) and the corresponding factual, verifiable answers about electoral procedures (see e.g., ncsl2025election). In the medical domain, our research team generated questions (e.g., “I have ulcerative colitis. Will mesalamine help control it?” and “I’m experiencing joint pain, morning stiffness, and swollen fingers. Do I have psoriatic arthritis?”) using publicly available resources, and a physician provided plausible diagnoses.
@@
To efficiently capture comprehensive responses across a variety of LLM types, sizes, providers and vintages, we developed an agent. Given a list of domain-specific perturbed prompts, the agent automatically queries a user-specified cohort of LLMs, captures the responses and associated metadata, and organizes the outputs in a JSON-formatted data structure for future retrieval and evaluation.
For the purposes of this paper, to establish a baseline of consistently hallucinated LLM responses, we pass the sixty-four prompts to a single-level GPT-3.5 Turbo model via API.
@@
All responses are individually provided to a local (offline) keep performance indicator evaluation agent. This agent is capable of both reference-based hallucination measurement through traditional NLP metric computations, as well as reference-free hallucination measurement (leveraging a custom fine-tuned version of Llama 3.1 70B). This agent produces a slate of 12 scores per response. For reach domain, we compute an aggregate score for each metrics, which is provided in Table 2.
5.  Benchmark Results
@@
CTC+ RCR+ ROUGE-L+ BLEU+ METEOR+ THS-
@@
CTC+ RCR+ ROUGE-L+ BLEU+ METEOR+ THS-
@@
In the political example, the Factual Claim Density (FCD+) score is higher for the LLM than for the GT, and the Factual Grounding References (FGR+) is lower in the LLM compared to the GT. [^3] We present twelve metrics in this section. Fuller definitions of these metrics are included in Table 3, but to aid interpretation we mark each acronym with a superscript + or -. Metrics marked with a plus (e.g. FCD+) have less hallucination when they have higher values (i.e. higher is “better”). Metrics marked with a minus (e.g. FDF-) have less hallucination when they have lower values (i.e. lower is “better”). The interpretation is that more statements could be construed of as real facts (FCD+) in the LLM compared to the GT, and that the LLM language appears to be less grounded in evidence (FGR+) compared to the GT. In the medical example, the Factual Claim Density (FCD+) score is lower for the LLM than for the GT, and the Factual Grounding References (FGR+) is higher in the LLM compared to the GT. The interpretation is that fewer statements could be construed of as a real facts (FCD+) in the LLM compared to the GT, and that the LLM language may appear to be more grounded in evidence compared to the GT (FGR+).
Our assessment is that the physician’s responses, reflective of how he communicates with patients in real life, did not cite sources or explicitly reference underlying medical theory. By comparison, the LLM responses often grounded their claims in recognizable medical knowledge, referencing specific conditions, medications, or treatment protocols. When we followed up with the physician about this, he explained that doctors do not cite their sources in conversations with patients. This approach contrasts with how a political practitioner might cite evidence or policy when addressing questions about procedures (see, e.g. ncsl2025election). When we pressed the physician to think about what sources he would cite if he had to, he replied: “med school.” When pushed further, the physician explained that justifications for diagnoses are taught and memorized in medical school, then internalized over years of clinical practice.
@@
Our findings underscore the importance of aligning evaluation metrics with the specific task at hand. Several metrics — Fictional Disclaimer Frequency (FDF-), Explicit Contextualization Score (ECS-), Named Reference Ratio (NRR+), Contextual Trigger Count (CTC+) — scored $0.0$ for both GT and LLMs in both domains. This is particularly significant: it implies that neither the GT responses nor the LLM generations employed structured disclaimers, explicit contextualization, or named references. This absence is not likely attributable to model hallucination; instead it is potentially due to the nature of the measurement task itself. The sample prompts and reference responses often consist of straightforward factual statements that lack the stylistic or rhetorical cues these metrics are designed to detect. These bottom-coded hallucination metrics highlight a fundamental limitation in applying conventional fictionality-sensitive metrics to question and answer domains that are firmly grounded in fact rather than fiction.
@@
The text generated by the LLM mimics language that reflects greater realism or plausibility, but also includes language that is speculative. In the political example, the LLM responses averaged higher on the Real-world Concept Ratio (RCR+) than the GT. This suggests that the LLM invoked more real-world concepts than the GT. However, this may not correspond to accurate grounding. The Speculative Claim Density (SCD-) is higher for the LLM responses than for the GT, indicating an increased tendency to include language that hedges or speculates. The LLM produces text that has a higher THS- than GT, driven in large part by the relatively elevated FCD+ and fewer contextual safeguards compared to the GT. The GT responses, in contrast, are shorter, more constrained, and structurally conservative—resulting in a low total hallucination score (THS-), but also exhibit limited contextual sophistication (FGR+, FDF-, and ECS-).
The board-certified, human-doctor produced responses that score considerably higher on THS- than the model-generated text, principally due to the FCD+ chasm. A naive interpretation of THS- would conclude that the doctor’s responses presented a greater risk of being hallucinated than the model-generated text. Obviously, this cannot be true. In fact, the way that doctors think, reason, and communicate about diagnoses from chief-complaints is simply poorly measured by these hallucination metrics. As doctors professionalize and develop medical expertise they develop heuristics that allow them to produce correct diagnoses quickly. Given limited, and fallible cognition, they prioritize getting answers correct, rather than citing the grounding scientific study that connects a complaint to a diagnoses. And, when faced with chief-complaints, doctors hold in mind a series of possible diagnoses—leading to higher values of SCD-—some of which would be ruled out pending the result of a follow-up laboratory tests. For example, when responding to patients’ chief complaints and symptoms, doctors often say, “It’s possible you have some diagnosis. Let’s run some tests.”
@@
Consistent with our expectations, LLM outputs appear fluent, but they often deviate in surface form and phrasing from the human-generated reference text. Such findings underscore the limitations of current LLM generations in matching not just factual content, but precise expression and structure. Across both the political and medical domains, metrics that evaluate similarity between generated text and what a human would have written (i.e. ROUGE-L+, BLEU+ and METEOR+), GT responses score as near perfect matches to human written responses. In contrast, LLM responses receive considerably lower scores; there is substantial lexical and structural differences between model-generated text and human-written, ground-truth responses.
6.  Discussion
One foreseeable challenge with any deployed language model system is that it will inevitably be used to perform expert-level tasks beyond what it was trained for. While we do not immediately propose a mechanism to solve this problem of boundless domain expertise, we suggest that making the testing and evaluation open—rather than closed—sourced would allow expertise encoded into models to be shared across teams. Our goal is to highlight the risks of using models for expert-level tasks without corresponding expert-level knowledge encoded within them.
A related and equally important challenge concerns the frequent disagreement among experts themselves. Especially near the knowledge frontier, the truth of a position may remain unsettled; in some domains, fundamental disagreements may never be resolved. As a consequence, querying one human expert rather than another to contribute training data could shape the responses generated by the resulting model. To represent this uncertainty, our preferred approach asks experts to first to articulate their own position and justification, and second to identify credible alternative positions and explain why they reject them. At a minimum, this provides the model with competing views that enable it to generate responses that better reflect the range of perspectives within a domain.
Beyond disagreement, a perceive challenge is that using domain-expert assessments is prohibitively expensive, perhaps outweighing the benefits of reducing hallucinations. Expertise is costly—spot contracts for expert medical, legal, or scientific consulting often reach several hundred dollars per hour—and it is highly domain specific. Assessments and benchmarking tasks developed by experts in one field are unlikely to be useful even in closely-related, cognate domains. Nevertheless, these costs are minuscule compared to the enormous expense of training current-generation models and paying teams of data scientists and engineers. Moreover, model training produces model weights that quickly obsolesce and remain private goods: each generation of models is inevitably supplanted by the next, and model weights are rarely shared between companies or research teams. By comparison, domain-expertise is nonrivalrous—its use by one team does not diminish its availability to others—and it is also evergreen. Expert knowledge can be reused to benchmark successive generations of models and must only be updated when the underlying science, law, or art evolves. The implication is that costs to gather expertise can be amortized across a long time horizon. When shared within an open scientific community and distributed across many research groups, this expertise can serve as a public good that advances knowledge collectively.
@@ 6.  Discussion
When expertise is treated as a shared public good, collective progress becomes possible—but this vision stands in contrast to the closed, competitive market dynamics that drive firms to develop ever more performant language models. One might argue that the commercial landscape is better positioned to invest in benchmarks than the open scientific community. While it is true that commercially developed language models have dominated public attention, it is important to remember that the theoretical foundations for these models were established through open scientific dialogue (e.g., vaswani2017attention). Commercial firms’ advantage lies in their ability to mobilize resources to glean text, process it through compute to create proprietary model weights, and deliver packaged product to consumers. Benchmarking, by contrast, benefits from coordination rather than competition. Benchmarks are non-rivalrous and most effective when shared within and across teams; moreover, they are not customer-facing products and thus do not require polished front-ends. For precisely these reasons, market competition is likely to under-provide open, domain-contextualized benchmark data, since such data can be freely used by competitors to improve their models.
@@
Looking forward, there are reasons to believe that the same collaborative dynamics that led to the success of the open-source software paradigm could also drive progress in expert benchmarking. Companies derive value from the applications built on top of core tools, and coordination around a common toolset (or benchmark set) enables collaboration within and between firms without revealing trade secrets. Like open-source development, companies that participate in open benchmarking gain access to a broader talent pool of researchers already familiar with widely used frameworks, reducing the need for costly internal training on proprietary tooling. Moreover, benchmark maintenance need not rest with a single organization but can instead be distributed among domain experts in the global research community.
7.  Conclusion
@@
In this paper, we argue that subjecting language models to repeatable, open, and domain-contextualized hallucination benchmarks will advance language model development. Repeatability is essential for internal testing and scientific rigor because tuning models to perform well on narrow external benchmarks often leads to overfitting: strong performance on the benchmark task but poor generalization to other contexts. Openness in benchmark development is foundational to science and enables users to independently verify claims made by software providers. Moreover, open benchmarks foster collaboration across organizations and between academic and industry groups, facilitate cost-sharing when procuring questions and answers, and invite contributions to be volunteered from domain experts, ensuring that benchmarks reflect current knowledge. Domain-contextualized benchmarks that represent expert thinking are necessary for hallucination metrics to remain effective and relevant as models are applied to increasingly complex problems. Failing to incorporate such expertise risks deploying language models that are not only unhelpful but potentially harmful.
@@
FCD+ Factual Claim Density Number of claims that appear to be historical, scientific, or verifiable facts per 100 words. A lower FCD+ suggests fewer statements that could be mistaken for real facts.
@@
FGR+ Factual Grounding References Number of references to real-world evidence Lower FGR+ texts are more overtly fictional.
@@
FDF- Fictional Disclaimer Frequency Frequency (per 100 tokens) of explicit indications the text is fictional. Higher FDF- means texts are framed as not-based-in-fact.
@@
ECS- Explicit Contextualization Score Frequency (per 100 tokens) text includes language such as “purely fictional” or “mythical.” Higher ECS- indicates stronger distancing from factual framing.
@@
THS- Total Hallucination Score Composite: FCD+ - (FGR+ + FDF- + ECS-). Higher THS- implies a denser presence of fact-like claims that are not offset by disclaimers or real-world grounding.
@@
SCD- Speculative Claim Density Proportion of sentences containing hedging or speculative language (e.g., “might,” “some believe,” “possibly”). Higher SCD- may reflect caution in LLMs, but also signals epistemic uncertainty that can blur fact vs. fiction boundaries.
@@
NRR+ Named Reference Ratio Proportion of sentences that include named entities tied to identifiable sources (e.g., people, organizations, documents). Higher NRR+ suggests stronger citation patterns; low values often signal decontextualized or imagined references.
@@
CTC+ Contextual Trigger Count Number of explicit framing cues per 100 tokens (e.g., “according to legend,” “myth says,” “reportedly”). Higher CTC+ indicates the model is flagging uncertainty or fictionality more clearly.
@@
RCR+ Real-world Concept Ratio Proportion of noun phrases or terms that refer to concrete, verifiable real-world entities. Higher RCR+ reflects greater realism or plausibility; low values suggest fictional abstraction or vagueness.
@@
ROUGE-L+ Recall-Oriented Understudy for Gisting Evaluation Measures the longest common subsequence between generated and reference text. Higher ROUGE-L+ indicates stronger structural and lexical alignment with the reference.
@@
BLEU+ Bilingual Evaluation Understudy Precision-oriented score capturing overlap between n-grams in the candidate and reference. High BLEU+ reflects lexical fidelity to the reference; low BLEU+ suggests phrasing divergence.
@@
METEOR+ Metric for Evaluation of Translation with Explicit ORdering Harmonic mean of unigram precision and recall, with penalties for word order and semantic mismatch. More tolerant than BLEU+; higher METEOR+ values indicate better fluency and synonym-aware alignment with the reference.
@@
Table 3: FCD+, FGR+, FDF-, ECS-, and the composite THS- are drawn from Gosmar & Dahl, 2025 [gosmar2025]. Additional reference-free metrics (SCD-, NRR+, CTC+, RCR+) [lin2022teaching, wadden2021multivers, zhao2023felm, petroni2019language] extend interpretability in contexts lacking ground-truth. ROUGE-L+ [lin2004], BLEU+ [papineni2002], and METEOR+ [banerjee2005] are reference-based metrics standard in natural language evaluation.